# Development and Benchmarking of JANGOFETT: A Novel Geant4-Operated Fission Event Tracking Tool

### v1

@@ 3.1 Dataset
Figure 4: Mass estimates based on KE from the 2E-method. The data are not unfolded or otherwise modified after the 2E loop is completed.
The test setup consisted of a system with two silicon detectors for measuring total kinetic energy (TKE) of fission fragments, surrounded by 21 HPGe detectors of varying geometries $\sim$20 cm from the source. Annular bismuth germanate (BGO) detectors surround the HPGEs for Compton suppression, which has been employed in the analysis used here.
Coincidence times between heavy fragments in silicon detectors are shown in Figure 2 and the coincidences between heavy fragments in the silicon detectors and gammas are shown within the supplemental material Figure S.2 , with a long tail that ends before 20 ns; beyond this point, false coincidences are about as likely as true. The average time for the two silicon detector signals was chosen as $\Delta_{time}=0$, and 50 ns coincidence windows (-30 ns to +20 ns) were chosen.
3.2 Kinetic Energy Simulations
Once coincidence windows were established, TKE was determined by summing the kinetic energies of coincident fission fragments. The resulting TKE distribution was fitted with a Gaussian, peaking at 180.083$\pm$0.001 MeV. Individual kinetic energy (KE) spectra were fitted with two Gaussians, with peaks at 79.544$\pm$0.002 and 102.229$\pm$0.001 MeV, which is within a reasonable calculational uncertainty of existing measurements for 252Cf as a calibration standard (181.03 MeV, 78.42 MEV, and 102.61) (Weissenberger et al., ). These are consistent with the CGMF outputs, indicating no bias in our analysis and steps-to-hits processing, other than being slightly low due to energy loss by X-rays and delta electrons.
Figure 3 shows a pair of KE (individual PFPs) and TKE (summed PFP energy) plots. Silicon detector signals with less than 20 MeV were excluded. As can be seen in Supplementary Figure S.3, there are a number of TKE signals that occur at $\sim$90 MeV and $\sim$180 MeV above the TKE peak - these are from false silicon detector triple and double coincidences that are present at a physically meaningful rate when considering the coincidence window. The TKE was analyzed before and after Geant4 simulations to rule out systematic errors -see the supplementary material for more detail (Figure S.4).
@@ 3.2 Kinetic Energy Simulations
The fission fragment kinetic energy was further used to approximate fragment mass via the double-energy (2E) method (Schmitt et al., ), yielding mass distributions with an accuracy of approximately ±4 atomic mass units (amu). Figure 4 demonstrates the 2E-method mass yields without any additional gating conditions beyond timing coincidence.
3.3 252Cf Fragment Identification
@@
With these coincidence criteria established, specific fission fragment pairs were selected for verification studies. Four heavy fragments were selected for gating, 144Ba, 141Cs, 149Nd, and 150Ce; the first is in the text of this article, while the remainder are included in the Supplementary Material. To isolate 144Ba (or 141Cs or 149149{149}149Nd or 150Ce) events, an initial mass gate was applied, selecting events within 4 amu of the desired nucleus, based on the 2E mass estimate. A second coincidence gate was then imposed, requiring $\gamma$-$\gamma$ coincidences from the heavy fragment to be measured in the HPGe detectors without a BGO Compton veto. $\gamma$-rays for coincidence were selected from a subset of high-probability $\gamma$-rays in the established nuclear data. Finally, an additional $\gamma$-gate was applied for each of the 0-5 neutron light fragment pairs. For these selected fragments, $\gamma$-$\gamma$-$\gamma$ spectroscopy histograms were generated to demonstrate both heavy and light fragments were simultaneously detected. It should be noted that gamma gating was not symmetric, e.g. if two gammas from the molybdenum isotope were identified but only one of the 144Ba, it was not included. Such a protocol would significantly increase the number of verified events, which was unnecessary to demonstrate the timing overlap of the PFPs, as was the goal here. Furthermore, no Doppler broadening was applied to energies for the purposes of analysis. Once applied, it is expected to improve statistics.
@@
All peaks in the histograms were evaluated to determine whether the identified fragments could be plausibly misidentified and should be discarded. Our evaluations did not determine that any fragments needed to be discarded in this way, as all gamma signals that were well above background were able to be identified with reasonable certainty as coming from the heavy or light fragment. The results from the photon coincidence plots are shown in Figure 5 (see Figure S.5 - S.7 for other gated fragments and Table 1-4 for a list of the $\gamma$ energies used to gate in the supplemental material).
(a) 0n
(b) 1n
(c) 2n
(d) 3n
(e) 4n
(f) 5n
[…]
@@
After identifying fragments from mass estimates and coincident $\gamma$, the KE of each fragment was summed to produced TKE plots for each of the light fragments observed, which are displayed in Figure 6.
(a) 0n
(b) 1n
(c) 2n
(d) 3n
(e) 4n
(f) 5n
[…]
@@ 4 Summary
JANGOFETT is a new tool for Geant4, developed to read primary fission product pairs into Geant4 with simulated time-based correlations. JANGOFETT performance was verified using a testbed of 252Cf(sf). Verification that the simulation produces distinguishable list-mode time sorted hits in detector volumes was performed by isolating coincidences from specific heavy and light fragments, based on mass estimates and photon energy. A heavy fragment of ${}^{1}44$Ba and its 0-5n light complements are shown here, with gamma-ray signals for each partner fragment appearing in the same simulation time, though they came from individual Geant4 instances. Additional fragments were analyzed and are included in the Supplementary Material. This tool enables the simulation and analysis of covariant fission observables in Geant4, generating data structures that are similar to those produced by experimental apparatus in fission for analysis.
@@ Acknowledgments
Authors LW and JS acknowledge fellowships from the Nuclear Regulatory Commission University Nuclear Leadership Program administered through Oregon State University. Authors gratefully acknowledge a startup package from the Oregon State University College of Engineering.

### v2

@@ 3.2 Kinetic Energy Response
The PFP KE was further used to approximate PFP mass via the double-energy (2E) method [21], yielding mass distribution with a precision of approximately $\pm$4 atomic mass units (amu). The pre-neutron evaporation FFs are then reconstructed via calculations of 252Cf(sf) with neutron multiplicity as a function of post-neutron evaporation from GEF[20]. This method has been utilized in similar mass yield measurements in recent articles [12][5][18]. Figure 4 demonstrates the 2E-method mass yields without any additional gating conditions beyond timing coincidence and compares estimated yields to input FF masses from CGMF in addition to yields produced from GEF.
Figure 4: Comparison of inferred FF mass distributions: (1) JANGOFETT 2E-method estimates based on KEs, without unfolding or further post-processing; (2) the input FF mass distribution from CGMF[22]; and (3) comparative predictions from the GEF code[20]. A version of the JANGOFETT distribution smoothed with a 5 amu-wide binning is also shown to illustrate overall yield trends and approximate experimental mass resolution[27]. Note the lower peaks and broader symmetric fission branch due to 2E method broadening, compared to the input from CGMF.
3.3 252Cf(sf) PFP Identification
@@
With these coincidence criteria established, specific PFP pairs were selected for verification studies. Four heavy PFPs were selected for gating, 141Cs, 144Ba, 150Ce 149Nd.144Ba is used for analysis in this work, while the remainder are included in the Supplementary Material. To isolate specific PFP events, an initial mass gate was applied, selecting events within 4 amu of the desired isotope, based on the 2E mass estimate. A second coincidence gate was then imposed, requiring a $\gamma$ coincidences from the heavy PFP to be measured in the HPGe detectors without a BGO Compton veto. $\gamma$-rays for coincidence were selected from a subset of high-probability $\gamma$-rays in the established nuclear data. Finally, an additional $\gamma$-gate was applied for each of the light PFP possible with 0-5 neutrons evaporated from the pair. In this analysis, only information that could be extracted from experimental methods was used to determine the PFP pair. As a result, the 0-5n values represent neutrons removed from the system agnostic of which FF they were removed from. For these selected PFPs, Coincident $\gamma$ histograms were generated to demonstrate both heavy and light PFP were simultaneously detected.
@@
All peaks in the histograms were evaluated to determine whether the identified fragments could be plausibly misidentified by mass and $\gamma$-gates and should be discarded. Our evaluations did not determine that any fragments needed to be discarded in this way, as all $\gamma$ signals that were above background were able to be identified with reasonable certainty. The results from the $\gamma$ singles plots when gated on 2E-masses are shown in Figure 5 (see Figure S.5 - S.7 for other gated fragments and Table S.1-S.4 for a list of the $\gamma$ energies used to gate in the Supplemental Material).
(a) 1n
(b) 2n
(c) 3n
(d) 4n
(e) 5n
(f) 6n
Figure 5: $\gamma$ energy spectra when at least one coincident $\gamma$ from 144Ba and at least one $\gamma$ from the correlated PFP (i.e. ${}^{252}Cf\rightarrow^{144}Ba+^{108-x}Mo+xn$), for $x$ between 1 and 6 for (a) through (f). Energy labels on the figure are given in keV with the respective PFP label.
@@
After identifying fragments from mass estimates and coincident $\gamma$, the KE of each fragment was summed to produced TKE plots for each of the light fragments observed, which are displayed in Figure 6. It is difficult to conclude much after only 24 hours of simulated data, but the presented figures are sufficient to conclude that multiple fragments are simulated within a single time-correlated event, distinguishable from random background.
(a) 1n
(b) 2n
(c) 3n
(d) 4n
(e) 5n
(f) 6n
Figure 6: TKE plots for tagged 144Ba + Light PFP events from at least one $\gamma$ from 144Ba and at least one $\gamma$ from the correlated light PFP. It should be noted that expected trends between TKE and TXE are not observed due to low statistics and overgeneralized $\gamma$ gates resulting in the inclusion of events not attributed to 144Ba + its Light PFP. This is expected to be improved with an increase in simulated fission events and more conservative gating requirements. This figure is included to demonstrate the ability to isolate multiple particle events using time correlations in Geant4 with JANGOFETT.
@@ 4 Summary
JANGOFETT is a new tool for Geant4, developed to read primary fission product pairs into Geant4 with simulated time-based correlations. JANGOFETT performance was verified using a simulation of 252Cf(sf) detector responses. Verification that the simulation produces distinguishable list-mode time sorted hits in detector volumes was performed by isolating coincidences from specific heavy and light PFPs, based on mass estimates and $\gamma$ energy. A heavy PFP of 144Ba and its 0-5n light complements are shown here, with $\gamma$ signals for each partner PFP appearing in the same simulation time, though they came from individual Geant4 instances. Additional PFPs were analyzed and are included in the Supplementary Material. This tool enables the simulation and analysis of covariant fission observables in Geant4, generating data structures that are similar to those produced by experimental apparatus in fission for analysis.
@@ Acknowledgments
Authors LW and JS acknowledge fellowships from the Nuclear Regulatory Commission University Nuclear Leadership Program administered through Oregon State University. Authors gratefully acknowledge a startup package from the Oregon State University College of Engineering. Author LW acknowledges additional support from the ARCS foundation Oregon.